From the following information from the payroll register of Veronica's Auto Supply Store, calculate the amount of taxable earnings for unemployment and FICA tax, and prepare the journal entry to record the employer's payroll taxes as of April 30, 20--. Social Security tax is 6.2% on the first $94,200 of earnings for each employee. Medicare tax is 1.45% of gross earnings. FUTA tax is 0.8%, and SUTA tax is 5.4% each on the first $7,000 of earnings.

Cumulative Pay

Current

Taxable Earnings


Employee Name

Before Current
Earnings

Gross
Pay

Unemployment
Compensation

Social
Security

Carnavale, Liz    $ 6,400

$1,800

O'Malley, Sean 6,850

1,800

Sarlo, Jarred 109,300

3,700

Mandel, Melissa 74,200

3,200

Davenport, Vin 27,500

1,800
Answer #1
Cumulative Pay Current Taxable Earnings
Employee Name Before Current
Earnings
Gross Pay Unemployment
Compensation
Social
Security
Carnavale, Liz 6400 1800 600 1800
O'Malley, Sean 6850 1800 150 1800
Sarlo, Jarred 109300 3700 0 0
Mandel, Melissa 74200 3200 0 3200
Davenport, Vin 27500 1800 0 1800
Total 12300 750 8600
Journal entry:
Debit Credit
Payroll tax expense 758.05
      Social Security tax payable 533.20 =8600*6.2%
      Medicare tax payable 178.35 =12300*1.45%
      FUTA tax payable 6.00 =750*0.8%
      SUTA tax payable 40.50 =750*5.4%
Workings:
Taxable earnings:
Unemployment Compensation:
Carnavale, Liz 600 =7000-6400
O'Malley, Sean 150 =7000-6850